OpenStudy (anonymous):

Henrys mp3 player has 512mb of memory he has already downloaded 287mbb and will continue to downloaded m more megabytes write and solve an inequality that shows how many more megabytes he can download

OpenStudy (anonymous):

it's unclear if we're looking for m or for the amount after he downloads m

OpenStudy (anonymous):

either: m = 512 - 287 = 225 or LEFT = 512 - 287 - m LEFT = 255 - m